(4x^3yz^4)(2xy^-2z^-1)(5x^-2y^-2z^3)
Answer by stanbon(48535)

(4x^3yz^4)(2xy^-2z^-1)(5x^-2y^-2z^3)
Multiply the coefficients: 4*2*5=40
Multiply the x-factors : x^3*x*x^-2=x
Multiply the y-factors : y^y^-2*y^-2=1/y^3
Multiply the z-factors : z^4*z^-1*z^3=z^6
----------
Multiply these results to get: 40xz^6/y^3
Cheers,
Stan H.